OpenStudy (anonymous):

The axis intercepts and the vertex of the curve y=x^2-12x+20 form the corners of a quadrilateral. I need to calculate the area..

OpenStudy (campbell_st):

i'd find the intercepts and vertex and determine the type of quadrilateral ok... so what are the intercepts y-interscept let x = 0 and solve for y x- intercepts let y = 0 and solve \[x^2 - 12x + 20 = 0\] the vertex find the line of symmetry x = -b/2a in your question b = -12 and a = 1 then substitute this value to into the equation to find the vertex... plot the points... then work from there...

OpenStudy (anonymous):

cheers!